# Multistable Shape from Shading Emerges from Patch Diffusion

### v1

@@ Abstract
Models for monocular shape reconstruction of surfaces with diffuse reflection—shape from shading—ought to produce distributions of outputs, because there are fundamental mathematical ambiguities of both continuous (e.g., bas-relief) and discrete (e.g., convex/concave) varieties which are also experienced by humans. Yet, the outputs of current models are limited to point estimates or tight distributions around single modes, which prevent them from capturing these effects. We introduce a model that reconstructs a multimodal distribution of shapes from a single shading image, which aligns with the human experience of multistable perception. We train a small denoising diffusion process to generate surface normal fields from $16\times 16$ patches of synthetic images of everyday 3D objects. We deploy this model patch-wise at multiple scales, with guidance from inter-patch shape consistency constraints. Despite its relatively small parameter count and predominantly bottom-up structure, we show that multistable shape explanations emerge from this model for “ambiguous” test images that humans experience as being multistable. At the same time, the model produces veridical shape estimates for object-like images that include distinctive occluding contours and appear less ambiguous. This may inspire new architectures for stochastic 3D shape perception that are more efficient and better aligned with human experience.
@@ 1 Introduction
From chiaroscuro in Renaissance paintings to the interplay of light and dark in Ansel Adams’ photographs, humans are masters at perceiving three-dimensional shape from variations of image intensity—shading—from a single image alone. Even though our visual experience is dominated by everyday objects, our perception of shape from shading has been shown to robustly generalize to synthetic “non-ecological” images invented by vision scientists. Some of these images have ambiguous (e.g., convex/concave) interpretations and lead to multistable perceptions, where one’s impression of 3D shape alternates between two or more competing explanations. Figure 1 shows an example adapted from [29], which is alternately interpreted as an indentation in a convex surrounding surface or a protrusion on a concave surrounding surface. Both explanations are physically correct because the same image can be generated from either shape under different lighting conditions.
@@
Figure 1: Many shapes (left) can explain the same image (middle) under different lighting, including flattened and tilted versions and convex/concave flips. The concave/convex flip in this example is also perceived by humans, often aided by rotating the image clockwise by 90 degrees. Previous methods for inferring surface normals (SIRFS [4], Derender3D [50], Wonder3D [31]) or depth (Marigold [24], Depth Anything [53]) produce a single shape estimate or a unimodal distribution. Our model produces a multimodal distribution that matches the perceived flip. (Image adapted from [29].)
@@
Instead, we approach monocular shape inference as a conditional generative process, and inspired by a long history of shape from shading with local patches [15, 22, 7, 27, 51, 17, 14], we present a bottom-up, patch-based diffusion model that can mimic multistable perception for ambiguous images of diffuse shading. Notably, our model is trained using images of familiar object-like shapes and has no prior experience with the ambiguous images that we use for testing. It is built on a small conditional diffusion process that is pre-trained to predict surface normals from $16\times 16$ image patches. When this patch process is applied in a non-overlapping manner with simple inter-patch shape consistency constraints, our model coordinates those patches in the sampling process and turns out to capture global ambiguities that are experienced by humans.
@@
An important attribute of our model is the way it handles lighting. It builds on the mathematical observation that shape perception can precede lighting inference [28]. It also adheres to the philosophy that inferred lighting cannot and should not be precise because of effects like global illumination [47]. Our model achieves this by guiding its diffusion sampling process with a very weak constraint on lighting consistency, where each patch nominates a dominant light direction and then all patches enact their own concave/convex flips in response to those nominations.
@@
Another critical aspect of our model is a diffusion sampling process that proceeds across multiple scales. It involves spatially resampling the normal predictions at intermediate diffusion time steps and then adding noise before resuming the diffusion at the resampled resolutions. Our approach is inspired by previous work that uses a “V-cycle” (fine-coarse-fine) to avoid undesirable local extrema during MRF optimization [34]. Our ablation experiments show that multi-scale sampling is crucial for finding good shape explanations that are globally consistent.
@@
We train our patch diffusion model on images of objects like those in Fig. 2(a), and we find experimentally that it can generalize to new objects as well as to images like Fig. 1, which are quite different from the training set and appear multistable to humans. This is in contrast to previous diffusion-based monocular shape models [24, 31], which cannot capture multistability and produce output that is much less diverse. Our model is also extremely efficient, only requiring a small pixel-based diffusion UNet that operates on $16\times 16$ patches. Our total model weights require only 10MB of storage, which is much less than the 2-3GB required by some of the previous models we compare to.
@@ 2 Background
2.1 Shape from Shading and its Ambiguities
Shape from shading is a classic reconstruction problem in computer vision. Since it was formulated by Horn in the 1970s [20], many approaches have been proposed to tackle it, often by assuming diffuse Lambertian shading and uniform lighting from either a single direction (e.g.,[40, 13]) or described by low-order spherical harmonics (e.g., [4, 56]). Almost all methods require the lighting to be known (e.g., for natural illumination [38]) or try to estimate it precisely via inverse rendering during the optimization process [51, 56, 50], and many approaches rely on a set of priors to constrain the possible search space. For instance, SIRFS [4] uses different lighting priors for natural versus laboratory conditions and a surface normal prior along occluding contours. Recent deep learning approaches like Derender3D [50] have demonstrated impressive results without being limited to Lambertian reflectance, but they similarly commit based on priors internalized from their training sets and have trouble generalizing to new conditions.
The main challenge of shape from shading comes from the many levels of inherent ambiguity. At a single Lambertian point, when lighting is unknown, there are multi-dimensional manifolds of surface orientations and curvatures that are consistent with the spatial derivatives of intensity at the point [27, 17]. Even when lighting is known, there is a cone of possible normal directions. At the level of a quadratic surface patch, when lighting is unknown there is generally a discrete four-way ambiguity corresponding to convex, concave and saddle shapes [51, 27]. At a global level, when lighting and surface albedo are unknown and perfectly uniform, the existence of continuous families of ambiguities arises from interpreting the Lambertian shading equation as a PDE without boundary conditions [26] or as a system of polynomial equations [11]. Moreover, any global shape explanation has a discrete counterpart that corresponds to a convex/concave flip. Finally, when the spatial albedo pattern is relaxed to allow for slight alterations, there is an additional three-parameter generalized bas-relief ambiguity [5] that corresponds to flattenings and tiltings of the global shape.
It is important to note that all of these mathematical ambiguities are based on certain idealized models for the image formation process, such as exact Lambertian shading, perfectly uniform albedo, and most commonly, perfectly uniform lighting that ignores global illumination effects such as interreflections and ambient occlusion which in reality have substantial effects [47]. An advantage of stochastic learning-based approaches like the one in this paper is that they have the potential to capture all of these ambiguities, as well as others that have not yet been discovered or characterized.
@@ 2.2 Denoising Diffusion with Guidance
Figure 2: Training patches are cropped from synthetic images of ordinary diffuse objects, and during training, a small diffusion model learns to denoise the normal field $x_{0}^{u}$ for patch $u$ from a random sample $x_{T}^{u}$ conditioned on the patch intensities $c^{u}$. During inference, the model is applied in parallel to non-overlapping patches, with guidance from inter-patch shape-consistency constraints.
Diffusion probabilistic models [19](DDPM) generate data by iteratively denoising samples from a Gaussian (or other) pre-determined distribution. We build on a conditional denoising diffusion model, where the condition $c$ is a grayscale image patch, and the model is designed to approximate the distribution $q(x_{0}|c)$ on 3-channel normal maps $x_{0}$ with a tractable model distribution $p_{\theta}(x_{0}|c)$. A ‘forward process’ adds Gaussian noise to a clean input $x_{0}$ and is modeled as a Markov chain with learned Gaussian transitions for timesteps $t=0,1,\cdots,T$. Each step in the forward process adds noise according to $q(x_{t}|x_{t-1},c):=\mathcal{N}(\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\mathbf{I})$, where $\{\beta_{t}\}$ is a predetermined noise variance schedule. The intermediate noisy input $x_{t}$ can be written as
@@
The ‘reverse process’ $q(x_{t-1}|x_{t},c)$ is modeled by a learned Gaussian transition $p_{\theta}(x_{t-1}|x_{t},c):=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t;c),\sigma_{t}^{2}\mathbf{I})$. The mean value $\mu_{\theta}(x_{t},t;c)$ can be expressed as a combination of the noisy image $x_{t}$ and a noise prediction $\epsilon_{\theta}(x_{t},t;c)$ from a learned model. The noise prediction model $\theta$ can be trained to minimize a prediction error $L({\theta}):=\mathbb{E}_{x_{0},\omega\sim\mathcal{N}(0,\mathbf{I})}\big{[}||\omega-\epsilon_{\theta}(x_{t},t;c)||_{2}^{2}]$ [19]. To sample noiseless data using the learned model, we start from an initial random Gaussian noise seed and use the learned denoiser $\epsilon_{\theta}$ to compute $x_{t-1}=\frac{1}{\sqrt{1-\beta_{t}}}(x_{t}-\frac{\beta_{t}}{\sqrt{1-\alpha_{t}}}\epsilon_{\theta}(x_{t},t;c))+\sigma_{t}z$ iteratively, where $z\sim\mathcal{N}(0,\mathbf{I})$, which is a stochastic process.
The denoising diffusion implicit model (DDIM) shows that the reverse procedure can be made deterministic by modeling it as a non-Markovian process with the same forward marginals [45]. This approach helps to accelerate the sampling process by using fewer steps and also provides an estimate of the predicted $\hat{x}_{0}$ at each timestep with $x_{t}$. Each denoising step combines noise and a forseen denoised version
@@
$f_{\theta}(x_{t},t;c)=\hat{x}_{0}(x_{t})=\frac{x_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}(x_{t};c)}{\sqrt{\alpha_{t}}}$ (3)
@@ 3 Multiscale Patch Diffusion with Guidance
Consider a differentiable surface with height function $h(x,y)$ over image domain $(x,y)$ viewed by parallel projection from above. The image plane is sampled on a square grid (pixels). We denote a $d\times d$ image patch as $c^{u}\in[0,1]^{d\times d}$ with patches indexed by $u$.
Training occurs on $d\times d$ patches extracted from images, as depicted in Fig. 2(a). For each patch $c^{u}$ there is a corresponding patch normal field $x_{0}^{u}\in[-1,1]^{3\times d\times d}$, whose $i$th spatial element represents a surface normal vector via
$\frac{x_{0}^{u}(i)}{\|x_{0}^{u}(i)\|}=\frac{(-p_{i},-q_{i},1)}{||(-p_{i},-q_{i},1)||}=n_{i}\in\mathbb{S}^{2},$ (6)
with $(p,q)=\left({\partial h}/{\partial x},{\partial h}/{\partial y}\right)$ the surface derivatives. Training proceeds as described in Sec. 2.2, with a dataset of patch tuples $(c^{u},x_{0}^{u})$ and a UNet $\epsilon_{\theta}(x^{u}_{t},t;c^{u})$ similar to that in [19], with four-channel input being the concatenation $c^{u}\oplus x_{t}^{u}$. (Model and training details are in Appendix A.2.)
As depicted in Fig. 2(b), inference occurs over $H\times W$ images $c$ that are divided into their collections of $d\times d$ non-overlapping patches $c^{u}$. We assume that the underlying global normal field $x_{0}\in[-1,1]^{3\times H\times W}$ is continuous at most locations, including at most locations on the seams between the patch fields $x^{u}_{0}$. We formulate the prediction of global field $x_{0}$ as reverse conditional diffusion on an undirected, four-connected graph $\mathcal{G}(\mathcal{V},\mathcal{E})$. Each patch $x_{0}^{u},u\in\mathcal{V}$ is a node, and there are edges between pairs of patches that are horizontally or vertically adjacent.
To encourage the patch fields to form a globally coherent prediction, we use guidance as described by Eqs. 4 and 5. Our guidance includes two terms:
@@
where $\mathcal{L}_{I}$ is a within-patch continuity term that encourages integrability of the normal fields over small pixel loops, and $\mathcal{L}_{S}$ is an inter-patch spatial consistency term that encourages constant curvature across the seams between patches. Hyperparameter $\lambda\in(0,1]$ controls the relative weighting of the two terms, and $\eta_{t}\in\mathbb{R}^{+}$ in Eq. 4 determines the overall guidance strength.
The integrability term follows Horn and Brooks [21] by penalizing deviation from a discrete approximation to the integrability of surface normals, i.e. ${\partial p}/{\partial y}={\partial q}/{\partial x}$, over $2\times 2$ loops of pixels. We write this as
@@
where the summation is over the $i,j$ grid-indexed pixels in patch $u$.

### v2

@@ Abstract
Models for inferring monocular shape of surfaces with diffuse reflection—shape from shading—ought to produce distributions of outputs, because there are fundamental mathematical ambiguities of both continuous (e.g., bas-relief) and discrete (e.g., convex/concave) types that are also experienced by humans. Yet, the outputs of current models are limited to point estimates or tight distributions around single modes, which prevent them from capturing these effects. We introduce a model that reconstructs a multimodal distribution of shapes from a single shading image, which aligns with the human experience of multistable perception. We train a small denoising diffusion process to generate surface normal fields from $16\times 16$ patches of synthetic images of everyday 3D objects. We deploy this model patch-wise at multiple scales, with guidance from inter-patch shape consistency constraints. Despite its relatively small parameter count and predominantly bottom-up structure, we show that multistable shape explanations emerge from this model for ambiguous test images that humans experience as being multistable. At the same time, the model produces veridical shape estimates for object-like images that include distinctive occluding contours and appear less ambiguous. This may inspire new architectures for stochastic 3D shape perception that are more efficient and better aligned with human experience.
@@ 1 Introduction
From chiaroscuro in Renaissance paintings to the interplay of light and dark in Ansel Adams’ photographs, humans are masters at perceiving three-dimensional shape from variations of image intensity—shading—from a single image alone. Even though our visual experience is dominated by everyday objects, our perception of shape from shading generalizes to many synthetic “non-ecological” images invented by vision scientists. Some of these images have ambiguous (e.g., convex/concave) interpretations and lead to multistable perceptions, where one’s impression of 3D shape alternates between two or more competing explanations. Figure 1 shows an example adapted from [31], which is alternately interpreted as an indentation or a protrusion. Both explanations are physically correct because the same image can be generated from either shape under different lighting conditions.
@@
Figure 1: Many shapes (left) can explain the same image (middle) under different lighting, including flattened and tilted versions and convex/concave flips. The concave/convex flip in this example is also perceived by humans, often aided by rotating the image clockwise by 90 degrees. Previous methods for inferring either surface normals (SIRFS [4], Derender3D [53], Wonder3D [33]) or depth (Marigold [27], Depth Anything [56]) produce a single shape estimate or a unimodal distribution. Ours produces a multimodal distribution that matches the perceived flip. (Image adapted from [31].)
@@
Instead, we approach monocular shape inference as a conditional generative process, and inspired by a long history of shape from shading with local patches [18, 25, 9, 29, 54, 20, 17], we present a bottom-up, patch-based diffusion model that can mimic multistable perception for ambiguous images of diffuse shading. Notably, our model is trained using images of familiar object-like shapes and has no prior experience with the ambiguous images that we use for testing. It is built on a small conditional diffusion process that is pre-trained to predict surface normals from $16\times 16$ image patches. When we apply this patch process at multiple scales with inter-patch shape consistency constraints, and when we coordinate the sampling across patches, the model ends up capturing global ambiguities that are very similar to those experienced by humans.
@@
An important attribute of our model is the way it handles lighting. It builds on the mathematical observation that shape perception can precede lighting inference [30]. It also adheres to the philosophy that inferred lighting cannot, and should not, be precise because of spatially-varying effects like global illumination [50]. Our model achieves these aims by guiding its diffusion sampling process with a very weak constraint on lighting consistency, where each patch nominates a dominant light direction and then all patches enact their own concave/convex flips in response to those nominations.
@@
Another critical aspect of our model is a diffusion sampling process that is coordinated across multiple scales. It involves spatially resampling the normal predictions at intermediate diffusion time steps and then adding noise before resuming the diffusion at the resampled resolutions. Our approach is inspired by previous work that uses a “V-cycle” (fine-coarse-fine) to avoid undesirable local extrema during MRF optimization [36]. Our ablation experiments show that multi-scale sampling is crucial for finding good shape explanations that are globally consistent.
@@
We train our patch diffusion model on images of objects like those in Fig. 2(a), and we find experimentally that it can generalize to new objects as well as to images like Fig. 1 which are quite different from the training set and appear multistable to humans. This is in contrast to previous diffusion-based monocular shape models [27, 33] which cannot capture multistability and produce output that is much less diverse. Our model is also extremely efficient, only requiring a small pixel-based diffusion UNet that operates on $16\times 16$ patches. Our total model weights require only 10MB of storage, much less than the 2-3GB required by some of the previous (and more general) models we compare to.
@@ 2 Background
2.1 Ambiguities in Shape from Shading
Shape from shading is a classic reconstruction problem in computer vision. Since being formulated by Horn in the 1970s [23] there have been many approaches to tackle it, often by assuming diffuse Lambertian shading and uniform lighting from either a single direction (e.g.,[43, 16]) or as low-order spherical harmonics (e.g., [4, 59]). Almost all methods either require the lighting to be known (e.g., for natural illumination [41]) or try to estimate it precisely via inverse rendering during the optimization process [54, 59, 53], and many approaches rely on a set of priors to constrain the possible search space. For instance, SIRFS [4] uses different lighting priors for natural versus laboratory conditions and a surface normal prior along occluding contours. Recent deep learning approaches like Derender3D [53] have demonstrated impressive results without being limited to Lambertian reflectance, but they similarly rely on priors internalized from their training sets and have trouble generalizing to new conditions.
The main challenge of shape from diffuse shading comes from the many levels of inherent ambiguity. At a single Lambertian point, when lighting is unknown, there is a multi-dimensional manifold of surface orientations and curvatures that are consistent with the spatial derivatives of intensity at the point [29, 20]. Even when lighting and surface albedo are known at a point, there is a cone of possible normal directions. At the level of a quadratic surface patch, when lighting is unknown, there is a discrete four-way ambiguity corresponding to convex, concave, and saddle shapes [54, 29]. At a global level, when lighting and surface albedo are known, ambiguities arise from interpreting the Lambertian shading equation as a PDE (e.g., [6]) or as a system of polynomial equations [13]. And when lighting and albedo are unknown, there is an additional three-parameter global ambiguity that corresponds to flattenings and tiltings of the global shape [5]. Finally, when lighting is unknown, a global shape has a discrete counterpart that corresponds to a global convex/concave flip.
It is important to note that all of these mathematical ambiguities are based on certain idealized models for the image formation process, such as exact Lambertian shading, perfectly uniform albedo, and most commonly, perfectly uniform lighting that ignores global illumination effects such as interreflections and ambient occlusion, which in reality have substantial effects [50]. An advantage of a stochastic, learning-based approach, like the one presented here, is the potential to capture all of these ambiguities as well as others that have not yet been discovered or characterized.
@@ 2.2 Denoising Diffusion with Guidance
Figure 2: Training patches are cropped from synthetic images of ordinary diffuse objects, and during training, a small diffusion model learns to denoise the normal field $x_{0}^{u}$ for patch $u$ from a random sample $x_{T}^{u}$ conditioned on the patch intensities $c^{u}$. During inference, the model is applied in parallel to non-overlapping patches, with guidance from inter-patch shape-consistency constraints to minimize the curvature smoothness loss $\mathcal{L}_{S}$ and integrability loss $\mathcal{L}_{I}$.
Diffusion probabilistic models [22](DDPM) generate data by iteratively denoising samples from a Gaussian (or other) pre-determined distribution. We build on a conditional denoising diffusion model, where the condition $c$ is a grayscale image patch, and the model is designed to approximate the distribution $q(x_{0}|c)$ on 3-channel normal maps $x_{0}$ with a tractable model distribution $p_{\theta}(x_{0}|c)$. A ‘forward process’ adds Gaussian noise to a clean input $x_{0}$ and is modeled as a Markov chain with Gaussian transitions for timesteps $t=0,1,\cdots,T$. Each step in the forward process adds noise according to $q(x_{t}|x_{t-1},c):=\mathcal{N}(\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\mathbf{I})$, where $\{\beta_{t}\}$ is a predetermined noise variance schedule. The intermediate noisy input $x_{t}$ can be written as
@@
The ‘reverse process’ $q(x_{t-1}|x_{t},c)$ is modeled by a learned Gaussian transition $p_{\theta}(x_{t-1}|x_{t},c):=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t;c),\sigma_{t}^{2}\mathbf{I})$. The mean value $\mu_{\theta}(x_{t},t;c)$ can be expressed as a combination of the noisy image $x_{t}$ and a noise prediction $\epsilon_{\theta}(x_{t},t;c)$ from a learned model. The noise prediction model $\theta$ can be trained by minimizing the prediction error
$L({\theta}):=\mathbb{E}_{x_{0},\omega\sim\mathcal{N}(0,\mathbf{I})}\big{[}||\omega-\epsilon_{\theta}(x_{t},t;c)||_{2}^{2}],$ (2)
as shown in [22]. To sample noiseless data using the learned model, we start from an initial random Gaussian noise seed and use the learned denoiser $\epsilon_{\theta}$ to compute $x_{t-1}=\frac{1}{\sqrt{1-\beta_{t}}}(x_{t}-\frac{\beta_{t}}{\sqrt{1-\alpha_{t}}}\epsilon_{\theta}(x_{t},t;c))+\sigma_{t}z$ iteratively, where $z\sim\mathcal{N}(0,\mathbf{I})$, which is a stochastic process.
The denoising diffusion implicit model (DDIM) shows that the reverse procedure can be made deterministic by modeling it as a non-Markovian process with the same forward marginals [48]. This approach helps to accelerate the sampling process by using fewer steps and also provides an estimate of the predicted $\hat{x}_{0}$ at each timestep with $x_{t}$. Each denoising step combines noise and a foreseen denoised version
@@
$f_{\theta}(x_{t},t;c)=\hat{x}_{0}(x_{t})=\frac{x_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}(x_{t},t;c)}{\sqrt{\alpha_{t}}}$ (4)
@@ 3 Multiscale Patch Diffusion with Guidance
Consider a surface represented by a differentiable height function $h(x,y)$ over image domain $(x,y)$ viewed by parallel projection from above. The image plane is sampled on a square grid (pixels). Image patches have size $d\times d$ and are indexed by $u$, and we denote them as $c^{u}\in[0,1]^{d\times d}$.
Training occurs on patches extracted from images of everyday objects, as depicted in Fig. 2(a). For each image patch $c^{u}$ there is a corresponding patch normal field $x_{0}^{u}\in[-1,1]^{3\times d\times d}$, whose $(i,j)$th spatial element represents a surface normal vector via
$\frac{x_{0}^{u}(i,j)}{\|x_{0}^{u}(i,j)\|}=\frac{(-p_{i,j},-q_{i,j},1)}{||(-p_{i,j},-q_{i,j},1)||}=n_{i,j}\in\mathbb{S}^{2},$ (7)
where $(p,q)=\left({\partial h}/{\partial x},{\partial h}/{\partial y}\right)$ are the surface derivatives. Training proceeds as described in Sec. 2.2, with a dataset of patch tuples $(c^{u},x_{0}^{u})$ and a UNet $\epsilon_{\theta}(x^{u}_{t},t;c^{u})$ similar to that in [22] whose four-channel input is the concatenation $c^{u}\oplus x_{t}^{u}$. (Model and training details are in Appendix A.2.)
As depicted in Fig. 2(b), inference occurs over images $c$ of size $H\times W$ which are divided into their collections of non-overlapping patches $c^{u}$. We assume that the underlying global normal field $x_{0}\in[-1,1]^{3\times H\times W}$ is continuous including at most locations on the seams between the patch normal fields $x^{u}_{0}$. We formulate the prediction of global field $x_{0}$ as reverse conditional diffusion on an undirected, four-connected graph $\mathcal{G}(\mathcal{V},\mathcal{E})$. Each patch $x_{0}^{u},u\in\mathcal{V}$ is a node, and there are edges $\{u,v\}\in\mathcal{E}$ between pairs of patches that are horizontally or vertically adjacent.
To encourage the patch fields to form a globally coherent prediction, we use guidance as described by Eqs. 5 and 6. Our guidance includes two terms:
@@
where $\mathcal{L}_{I}$ is a within-patch continuity term that encourages integrability of the normal fields over small pixel loops, and $\mathcal{L}_{S}$ is an inter-patch spatial consistency term that encourages constant curvature across the seams between patches. Hyperparameter $\lambda\in(0,1]$ controls the relative weighting of the two terms, and $\eta_{t}\geq 0$ in Eq. 5 determines the overall guidance strength.
The integrability term follows Horn and Brooks [24] by penalizing deviation from a discrete approximation to the integrability of surface normals, i.e., ${\partial p}/{\partial y}={\partial q}/{\partial x}$, over $2\times 2$ loops of pixels. We write this as
@@
where the summation is over the $i,j$ grid-indexed pixels in patch $u$, and $p,q$ are the components of $\hat{x}_{0}^{u}$ implied by Eq. 7.